OpenStudy (winsteria):

What is the solution to the equation? 4(3 + 7) = 4 · x + 4 · 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

OpenStudy (winsteria):

@Zerrena27

OpenStudy (danjs):

4(10)=4x + 28

OpenStudy (danjs):

40-28 = 4x

OpenStudy (danjs):

12 = 4x x=3

OpenStudy (winsteria):

So A?

OpenStudy (danjs):

well test it by putting 3 in for x and see if both sides are equal

OpenStudy (winsteria):

yay it is can you help me with another one?

OpenStudy (danjs):

sure

OpenStudy (anonymous):

Me as well.

OpenStudy (winsteria):

Okay

OpenStudy (winsteria):

Use the distributive property to rewrite the expression and solve. 9 ∗ 72 = 9 ∗ [?] + 9 ∗ 2 = [?]

OpenStudy (winsteria):

reload if there's the question mark glitch oops

OpenStudy (danjs):

ok

OpenStudy (anonymous):

Ok, this a hard one.

OpenStudy (anonymous):

x=3.

OpenStudy (danjs):

648 = 9x +18

OpenStudy (danjs):

i think you wrote it wrong

OpenStudy (anonymous):

yeah would it be 4?

OpenStudy (danjs):

rewrite using parethesis

OpenStudy (winsteria):

Here I took a picture of the question

OpenStudy (anonymous):

@Zerrena27 no its not. You just plug in the numbers and solve. I got 3

OpenStudy (danjs):

is it 9*72 = (9*x)+(9*2)

OpenStudy (anonymous):

Oh, I'm so dumb.

OpenStudy (danjs):

so 9*72 = 9(x+2)

OpenStudy (anonymous):

I didn't say that Zerrena

OpenStudy (winsteria):

Here's the question in picture form maybe you'll understand it better? http://oi57.tinypic.com/m8l07k.jpg

OpenStudy (anonymous):

I know, but i'm getting wrong answers.

OpenStudy (danjs):

i c

OpenStudy (anonymous):

me too, i think I got it.

OpenStudy (danjs):

9*72 = 9*36 + 9*2 = 648

OpenStudy (anonymous):

Yes!

OpenStudy (danjs):

they just want you to factor the 72 into 36*2

OpenStudy (anonymous):

Danjs you got a fan!

OpenStudy (winsteria):

So in the first box I put 36 and in the second one I put 648?

OpenStudy (danjs):

yea

OpenStudy (winsteria):

Thank you so much I HAVE ONE MORE CAN YOU HELP?

OpenStudy (anonymous):

Yea

OpenStudy (danjs):

sure

OpenStudy (anonymous):

what do medals do?

OpenStudy (winsteria):

They're just like awards?

OpenStudy (danjs):

not sure, i just started this site yesterday

OpenStudy (danjs):

just here for fun. hah

OpenStudy (anonymous):

I started today

OpenStudy (winsteria):

Here's my other question http://oi57.tinypic.com/bfqkno.jpg

OpenStudy (anonymous):

Whoa, what grade are you in?

OpenStudy (danjs):

12.2 239.12

OpenStudy (anonymous):

Is this area or perimeter?

OpenStudy (danjs):

me?

OpenStudy (anonymous):

yea

OpenStudy (danjs):

oh, 4th year mechanical engineering. lol

OpenStudy (winsteria):

So do you know what the answer is?

OpenStudy (danjs):

i put them both above

OpenStudy (danjs):

That problem was just factoring out the 2

OpenStudy (danjs):

to rewrite the expression , then just calculate it

OpenStudy (danjs):

12.2 239.12

OpenStudy (winsteria):

so the first box is 12.2 the second box is 239.12?

OpenStudy (danjs):

yes

OpenStudy (danjs):

wait

OpenStudy (winsteria):

The seconds box won't let me put in 239.12? it says its too big of a number

OpenStudy (danjs):

44

OpenStudy (danjs):

second box is 44

OpenStudy (danjs):

2*(12.2+9.8) = 2*22 = 44

OpenStudy (winsteria):

the first one is still 12.2 right?

OpenStudy (danjs):

yes P = 2*l + 2*w = 2(l+w)

OpenStudy (danjs):

it is just factoring out a 2

OpenStudy (winsteria):

okay and can you check this other answer of mine I just want to make sure it's right?

OpenStudy (danjs):

ok

OpenStudy (winsteria):

http://oi60.tinypic.com/apbcwp.jpg

OpenStudy (winsteria):

I think it's A

OpenStudy (winsteria):

oh wait did you already help me with that one o_o

OpenStudy (danjs):

4(3+7) = 4*x + 4*7 Using the same process as the last problem, factor out a 4 from the 4(3+7)

OpenStudy (danjs):

= 4*3+4*7

OpenStudy (winsteria):

is my answer right or no?

OpenStudy (danjs):

i mean distribute the 4 into the parenthesis

OpenStudy (winsteria):

ok

OpenStudy (danjs):

right a*(b+c) = a*b + a*c

OpenStudy (winsteria):

ok I'm right :)

OpenStudy (winsteria):

I need help with checking another one

OpenStudy (danjs):

ha ok one more, then open a new thread

OpenStudy (winsteria):

okay